OpenStudy (anonymous):

please help Select the ordered pair from the choices below that satisfies following system of equations: 2x - 10y = -14 4y = -x - 5 (-1/2, 1) (-20, 3) (3, 2) (8, -4)

jimthompson5910 (jim_thompson5910):

Simply check each answer choice

jimthompson5910 (jim_thompson5910):

So let's check answer choice A) 2x - 10y = -14 2(-1/2) - 10(1) = -14 -2/2 - 10 = -14 -1 - 10 = -14 -11 = -14 which is false So (-1/2, 1) does NOT lie on the line 2x-10y = -14 So choice A is NOT the answer

jimthompson5910 (jim_thompson5910):

go through the rest of the answer choices and check them keep in mind that the ordered pair must work for BOTH equations